OpenStudy (anonymous):

Jamie had the following problem on her math test: The product of 4 and a number is 12 more than 92. What is the number? When Jamie solved the problem, she wrote the following: 4n + 12 = 92 4n + 12 - 12 = 92 - 12 4n = 80 4n/4 = 80/4 n = 20 Then number is 20. A. Jamie's solution is not correct. Using complete sentences, explain what Jamie did wrong. (5 points) B. Write the correct algebraic equation for the problem that Jamie was to solve. (5 points) C. Solve the equation to find the number. Remember to use algebraic transformations. (5 points)

OpenStudy (karategirl2002):

First, let’s solve the problem. The product of 4 and a number is 12 more than 92. What is the number? It will be easier to see what is has been asked if you rewrite the sentence as follows: 4 times a number is 12 more than 92. What is the number? Substitute a multiplication sign (*) for “times”, substitute the variable n for “a number”, substitute an equal sign (=) for “is”, and substitute a plus sign (+) for “more than”. 4*n = 12 + 92 (this could also be written: 4*n = 92 + 12) 4n = 104 4n/4 = 104/4 n = 26 The number is 26 Using complete sentences, describe the student error. The student placed the +12 on the left side of the equation instead of the right side. The +12 should be placed on the right side of the equation

OpenStudy (anonymous):

@karategirl2002 can u number it lke A then B Then C pleez

OpenStudy (anonymous):

@karategirl2002

OpenStudy (karategirl2002):

Jamie did not interpret the words into an equation correctly. 4n=12+92 4n=104 n=26

OpenStudy (anonymous):

so that's A @karategirl2002

OpenStudy (anonymous):

@karategirl2002

OpenStudy (anonymous):

@karategirl2002

OpenStudy (karategirl2002):

yeah